OpenStudy (anonymous):

1. The volume of a cube is 8 cm^3. What is the surface area of this cube? 2 cm^2 4 cm^2 16 cm^2 24 cm^2

OpenStudy (anonymous):

24

OpenStudy (kinggeorge):

Well, we know the side length is 2 since \(2= \sqrt[3]{8} \) so each side has area \(2^2 =4\) so the total surface area should be \(4*6=24\) since a cube has 6 sides.